OpenStudy (anonymous):

Find the Foci for this Hyperbola: 16x^2-16y^2+128x-96y+368=0

OpenStudy (loser66):

first complete the square, can you do it?

OpenStudy (anonymous):

No that's why I'm asking you guys

OpenStudy (loser66):

I want to confirm the last number, it's +368 or -368?

OpenStudy (anonymous):

+368

OpenStudy (loser66):

ok divide both sides by 16, what do you have?

OpenStudy (anonymous):

0 and 23?

OpenStudy (loser66):

nope, divide all terms, like 16x^2/16 =x^2 16y^2/16 = y^2 and so on.

OpenStudy (anonymous):

oh ok

OpenStudy (loser66):

get this part?